decide whether or not the ordered pair is a solution to the equation 4x+2=24; (4,4). show all work to get the answer
Answer by FrankM(1040)

4x+2=24
4x=22
x=5.5
the pair is not a solution.
alternate answer -
4(4)+2=24
16+2=24
18=24
no. the 4 we entered into the equation produces an invalid result.